Good Time Company is a regional chain department store. It will remain in business for one more year. The probability of a boom year is 60 percent and the probability of a recession is 40 percent. It is projected that the company will generate a total cash flow of $205 million in a boom year and $96 million in a recession. The company's required debt payment at the end of the year is $130 million. The market value of the company’s outstanding debt is $103 million. The company pays no taxes.

  

a.

What payoff do bondholders expect to receive in the event of a recession? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)

b. What is the promised return on the company's debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What is the expected return on the company's debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a). Expected payoff to the bondholders is the lesser amount of the face value of debt or the company's value.

In the case of recession, the value of the company is $96 million, whereas the face value of debt is $130 million. So, the expected payoff to the bondholders is $96 million, the lesser of the two amounts.

b). Promised Return = [Face Value / Market Value of Debt] - 1

= [$130 million / $103 million] - 1

= 1.2621 - 1 = 0.2621, or 26.21%

c). In part a, we have established that the bondholders will receive $96 million in the case of recession. In case of boom, the bondholder will receive the full face value of $130 million as the value of company in that case is more than the face value.

Expected Payment to bondholders = [Prob.(Boom) * Payoff(Boom)] + [Prob.(Recession) * Payoff(Recession)]

= [0.6 * $130 million] + [0.4 * $96 million] = $78 million + $38.4 million = $116.4 million

Expected Return = [Expected Value / Market Value of Debt] - 1

= [$116.4 million / $103 million] - 1

= 1.1301 - 1 = 0.1301, or 13.01%